# Test-Time Canonicalization by Foundation Models for Robust Perception

### v1

@@ A.9 Comparison against Color Correction baselines
We benchmark our method on the RCC dataset and compare it against Barron & Tsai (2017) and the gray world method. We search for the two log-chrominance values in the range $[-0.7,-0.3]$, and used both classifier and diffusion energies with the default values for classifier energy, and a coefficient of $-10$ for diffusion energy, with the diffusion energy evaluated on a single step $t=50$. We used BO to optimize the color using $10$ random initial points and $50$ BO iterations. Our method only produces visually typical (in-distribution) images, not necessarily the most color-neutral version, and thus, performs poorly on this test compared to baselines designed to achieve high color accuracy on this task (Figure 17):. Specifically, we achieve a median angular error of $6.4^{\circ}$ compared to Barron & Tsai (2017)’s $1.3^{\circ}$, though we do beat gray-world’s $9.97^{\circ}$. Our method is nowhere near the specialized state-of-the-art methods on this task, but ours is a more general algorithm designed to improve invariance. We leave it to future work to improve the base performance of FoCal against supervised color approaches.
@@ A.10 Computational Complexity Analysis
[…]
Assuming image rotation uses negligible FLOPs, we get: 1×1$\times$1 × for naive, 8$\approx$ 8$\times$ 8 × for TTA, 56$\approx$ 56$\times$ 56 × for ours.
@@
[…]
@@ B.1 Experimental Setup - 3D
For Objaverse-LVIS (Deitke et al., 2023) in the “vary” stage we sample 60 viewpoints. These are at every 30 degrees of elevation and azimuth, for a grid of $[-60,-30,0,30,60]\times[-180,-150,-120,-90,-60,-30,0,30,60,90,120,150,180]$. For CO3D (Reizenstein et al., 2021), we sample the 12 viewpoints at the elevation 30 and azimuths of $[-180,-150,-120,-90,-60,-30,0,30,60,90,120,150,180]$.
@@ B.2 Experimental Setup – Color and Contrast
where $z=\sqrt{\exp(-L_{u})^{2}+\exp(-L_{v}^{2})+1}$ is a normalizing constant and $L_{u},L_{v}$ are the log-chroma values sampled from the uniform distribution with range $[-1,1]$. Intuitively, the log-chroma space defines the $R/G$ and $B/G$ ratios in log-space. A range of $[-1,1]$ corresponds roughly to a 7×7$\times$7 × change in the ratio between the minimum and maximum points of the range.

### v2

@@ A.9 Comparison against Color Correction baselines
We benchmark our method on the RCC dataset and compare it against Barron & Tsai (2017) and the gray world method. We search for the two log-chrominance values in the range [0.7,0.3][-0.7,-0.3], and used both classifier and diffusion energies with the default values for classifier energy, and a coefficient of 10-10 for diffusion energy, with the diffusion energy evaluated on a single step $t=50$. We used BO to optimize the color using $10$ random initial points and $50$ BO iterations. Our method only produces visually typical (in-distribution) images, not necessarily the most color-neutral version, and thus, performs poorly on this test compared to baselines designed to achieve high color accuracy on this task (Figure 17):. Specifically, we achieve a median angular error of $6.4^{\circ}$ compared to Barron & Tsai (2017)’s $1.3^{\circ}$, though we do beat gray-world’s $9.97^{\circ}$. Our method is nowhere near the specialized state-of-the-art methods on this task, but ours is a more general algorithm designed to improve invariance. We leave it to future work to improve the base performance of FoCal against supervised color approaches.
@@ A.10 Computational Complexity Analysis
$\text{Standard inference}:$ $\quad 1\times\text{(CLIP)}$
$\text{TTA}:$ $\quad 8\times\left(\text{Rotation}+\text{CLIP}\right)$
$\text{Our method}:$ $\quad 8\times\left(\text{Rotation}+\text{CLIP}+5\times\text{Diffusion}\right)$
Assuming image rotation uses negligible FLOPs, we get: $1\times$ for naive, $\approx 8\times$ for TTA, $\approx 56\times$ for ours.
@@
$\text{Theoretical Latency (parallel)}=\max\left(\text{Classifier},\text{Diffusion step},\text{CLIP}\right)$
@@ B.1 Experimental Setup - 3D
For Objaverse-LVIS (Deitke et al., 2023) in the “vary” stage we sample 60 viewpoints. These are at every 30 degrees of elevation and azimuth, for a grid of $60,-30,0,30,60]\times[-180,-150,-120,-90,-60,-30,0,30,60,90,120,150,180]$. For CO3D (Reizenstein et al., 2021), we sample the 12 viewpoints at the elevation 30 and azimuths of [180,150,120,90,60,30,0,30,60,90,120,150,180][-180,-150,-120,-90,-60,-30,0,30,60,90,120,150,180].
@@ B.2 Experimental Setup – Color and Contrast
where $z=\sqrt{\exp(-L_{u})^{2}+\exp(-L_{v}^{2})+1}$ is a normalizing constant and $L_{u},L_{v}$ are the log-chroma values sampled from the uniform distribution with range [1,1][-1,1]. Intuitively, the log-chroma space defines the $R/G$ and $B/G$ ratios in log-space. A range of [1,1][-1,1] corresponds roughly to a $7\times$ change in the ratio between the minimum and maximum points of the range.